OpenStudy (anonymous):

True or False? Y varies directly with x and y = -9 when x = 3. When y = 27, the value of x is 9.

OpenStudy (anonymous):

y = k/x 12 = k/6 12 X 6 = k 72 = k The value of the variation constant k is 72. so,,, yes